### v1

# Manuscript

Manuscript
Version One
ON THE REALIZABILITY OF
PRIME CONJECTURES
IN HEYTING ARITHMETIC
Milan Rosko
(November 2025)
We show that no total functional can uniformly transform $\Pi_{1}$ primality into explicit $\Sigma_{1}$ witnesses without violating normalization in $\mathsf{HA}$. The argument proceeds through three complementary translations: a geometric interpretation in which compositeness and primality correspond to local and global packing configurations; a proof-theoretic analysis demonstrating the impossibility of uniform $\Sigma_{1}$ extraction; and a recursion-theoretic formulation linking these constraints to the absence of total Skolem functions in $\mathsf{PA}$. The formal analysis in constructive logic is followed by heuristic remarks interpreting the results in informational terms.
Keywords: Proof Theory, Constructive Logic, Realizability;
Heyting Arithmetic; Goldbach’s Conjecture; Geometric Logic
MSC: 03F55, 03F03, 11A41, 68Q05, 11U05
Milan Rosko is from University of Hagen, Germany
Email: Q1012878 @@ studium.fernuni-hagen.de
ORCID: 0009-0003-1363-7158
Licensed under “Deed” ΓΔ  – creativecommons.org/licenses/by/4.0
1 Introduction
Let mathematics correspond to the greatest rigor after assumption; then logic will be the rigor before assumption. We examine the boundary where the two intersect. Classical arithmetic proceeds as though its objects—numbers, primes, and sums—are already well formed, whereas logic asks whether the rules that make such objects definable can themselves support constructive realization. The contrast becomes evident in statements such as Goldbach’s:
“Every integer that can be written as the sum of two primes can also be written as the sum of as many primes (including unity) as one wishes, until all terms are units.” Goldbach [1742], in a letter to Leonhard Euler.
Our analysis shows that any uniform construction transforming an even input into explicit prime summands would conflict with the normalization theorems of constructive proof theory rather than with any property of the primes themselves. It concerns the structure of realizability in arithmetic, not the distribution of prime numbers.
Figure 1: Predicting the decomposition of cuboids would effectively provide an oracle. Can the semantics of primality (gold) be syntactically separated from the “combinatorial entropy” of decompositions (right)? This question frames the present investigation.
The distinction between additive and multiplicative reasoning marks a precise logical frontier. To demonstrate “compositeness” is to exhibit explicit factors, a $\Sigma_{1}$ act. To demonstrate “primality” is to assert the universal absence of nontrivial factors, a $\Pi_{1}$ claim.
Roadmap. Section 2 establishes the constructive framework, defining the logical and semantic background within HA and the BHK interpretation of “proofs as realizers.” Section 3 develops the geometric translation of arithmetic predicates, where compositeness and primality are represented as finite configurations, exposing the $\Sigma_{1}/\Pi_{1}$ asymmetry in local versus global definability. Section 4 derives the corresponding logical barriers within HA. Section 5 reformulates these limits as inductive and fixed-point reductions. Section 6 translates the constructive constraints into classical recursion-theoretic and proof-length results within PA, including the oracle and speedup implications. Section 7 interprets the constructive and classical findings heuristically. Section 8 concludes with the epistemic and computational implications of these results, identifying the boundary between arithmetic truth and constructive realizability.
2 Preliminaries
2.1 Semantics
Definition 2.1 (Base Theory).
We consider first-order Heyting Arithmetic ($\mathsf{HA}$); see Heyting [1930]; Troelstra [1973]; Beeson [1985]; Buss [1998]; Hájek and Pudlák [1998] for details on $\mathsf{PA}$’s strictly weaker intuitionistic counterpart. Our language
$\mathcal{L}_{A}=\{0,S,+,\cdot,<\}$ (2.1)
is governed by constructive logic and includes full induction for all formulas of $\mathcal{L}_{A}$. Connectives are interpreted under the BHK semantics, following Brouwer [1975]; Heyting [1930]; Kolmogorov [1925]. In brief, the meaning of a formula is given by the construction that realizes it.
Definition 2.2 (Proof Semantics).
Proofs correspond to explicit constructions (realizers):
[…]
Thus, a proof in HA denotes a total, strongly normalizing operation transforming realizers into realizers; see Troelstra and van Dalen [1988]; Beeson [1985].
2.2 Thesis
Example 2.3.
Recall the Arithmetic Hierarchy of Kleene [1943]. It classifies subsets of $\mathbb{N}$ and arithmetical formulas according to the alternation of quantifiers in their definitions.
  1. 1.
    $\Delta^{0}_{0}$ (also $\Sigma^{0}_{0}=\Pi^{0}_{0}$): formulas containing only bounded quantifiers.
    $\varphi(x)\;:\;\exists y\leq x\;(y^{2}=x)$ (2.3)
    expresses $x$ is a perfect square” and is $\Delta^{0}_{0}$.
  2. 2.
    $\Sigma^{0}_{1}$: formulas beginning with an unbounded existential quantifier followed by a bounded (decidable) predicate.
    $\varphi(x)\;:\;\exists y\;T(x,y)$ (2.4)
    where $T(x,y)$ is a decidable relation (e.g., $y$ encodes a halting computation of program $x$). Example: “Program $x$ halts” is $\Sigma^{0}_{1}$.
  3. 3.
    $\Pi^{0}_{1}$: universal closure of a decidable predicate.
    $\psi(x)\;:\;\forall y\;P(x,y)$ (2.5)
    with $P$ decidable. Examples: “Program $x$ never halts” or “The sequence generated by $x$ is total” are $\Pi^{0}_{1}$.
  4. 4.
    $\Sigma^{0}_{n+1}$ and $\Pi^{0}_{n+1}$ are obtained by quantifier alternation:
    $\Sigma^{0}_{n+1}:\exists x_{1}\forall x_{2}\exists x_{3}\cdots\varphi_{0}(x_{1},\ldots,x_{n}),$ (2.6)
    $\Pi^{0}_{n+1}:\forall x_{1}\exists x_{2}\forall x_{3}\cdots\varphi_{0}(x_{1},\ldots,x_{n}),$
    where $\varphi_{0}$ is bounded ($\Delta^{0}_{0}$).
Thesis 2.4 (Realizability Barrier).
No total primitive recursive functional $E$ can uniformly transform $\Pi_{1}$ predicates—i.e., elements of $\operatorname{Prime_{\Pi}}$—into explicit $\Sigma_{1}$ witnesses within HA. Equivalently, no uniform $\Sigma_{1}$ totalizer exists for any nontrivial or combinatorial $\Pi_{1}$ property, as shown by Kleene [1952]; Friedman [1975]. This constraint remains invariant under any external extension, owing to syntactic parsimony.
2.3 Operators
Definition 2.5 (Basic Predicates).
Let $\mathbb{N}$ denote the natural numbers with successor $S$, addition, and multiplication. The following relations are definable in the language of HA:
[…]
Here, $\operatorname{Even}(n)$, $\operatorname{Odd}(n)$, and $\mathsf{Comp}(n)$ are $\Sigma_{1}$, while $\operatorname{Prime_{\Pi}}(n)$ is $\Pi_{1}$. This classification reflects the constructive asymmetry between witness-producing and universally quantified statements.
Remark.
Under Markov’s Principle, $\neg\neg\mathsf{Comp}(n)\to\mathsf{Comp}(n)$, but this principle is not derivable in HA. Consequently, $\operatorname{Prime_{\Pi}}(n)$ is not constructively decidable, even though it is arithmetically $\Pi^{0}_{1}$ in PA.
2.4 Constructive Translation
Definition 2.6 (Constructive Goldbach’s Principle).
Consider a constructive variant of Goldbach’s Binary Conjecture, The Strong Goldbach’s Conjecture:
$\mathrm{GC}^{*}_{\Pi}:\quad\forall n\Big(\operatorname{Even}(n)\Rightarrow\exists p,q\leq n\,[\operatorname{Prime_{\Pi}}(p)\wedge\operatorname{Prime_{\Pi}}(q)\wedge p+q=n]\Big).$ (2.8)
This statement combines $\Sigma_{1}$ witnesses $(p,q)$ with $\Pi_{1}$ side-conditions $(\operatorname{Prime_{\Pi}}(p),\operatorname{Prime_{\Pi}}(q))$. A constructive proof of (2.8) in HA would yield, via the Curry–Howard Correspondence, a total primitive-recursive operator
$E:\mathbb{N}\longrightarrow\mathbb{N}^{5},\quad E(n)=\langle p,q,a,b,c\rangle,$ (2.9)
where $a,b$ realize the predicates $\operatorname{Prime_{\Pi}}(p),\operatorname{Prime_{\Pi}}(q)$, and $c$ realizes $p+q=n$. Such an operator would act as a uniform $\Sigma_{1}$ extractor for $\Pi_{1}$ evidence, contradicting the strong normalization of HA proofs, see [Kreisel, 1951].
3 Geometric Logic
3.1 Topological Semantics
Figure 2: A $2\times 2$ rectangular packing realizing $\mathsf{Comp}(4)$. The configuration makes the multiplicative structure of 44 explicit (blue) within the grid semantics of $\mathcal{G}$.
Exposition.
Classical Gödel Completeness—that every consistent first-order theory has a model, [Gödel, 1930]—is not provable according to BHK, since it requires nonconstructive principles, e.g., Compactness. However, a constructive analogue holds for “geometric logic”, whose semantics are stable under “pullback” along morphisms.
Lemma 3.1 (Constructive Completeness).
Let $\mathcal{G}$ be a geometric theory and $\sigma$ a sequent. By [Fourman and Scott, 1979],
$\mathcal{G}\vdash\sigma\quad\Longleftrightarrow\quad\forall\mathcal{E}\in\mathsf{Topos},\;\forall\mathcal{M}\in\mathrm{Mod}_{\mathcal{E}}(\mathcal{G}),\;\mathcal{M}\models\sigma.$ (3.10)
In particular, taking $\mathcal{E}=\mathbf{Set}$ gives soundness and completeness with respect to ordinary (set-based) models.
Lemma 3.2 (Geometric Reduction Principle).
For any geometric theory $\mathcal{G}$ and geometric sequent $\sigma$,
$\models_{\mathrm{G}}\sigma\quad\Longleftrightarrow\quad\models_{\mathsf{L}}\sigma$ (3.11)
for every logic $\mathsf{L}$ whose semantics are reducible to geometric semantics.
Exposition.
Hence, “truth in geometric logic” coincides with “truth in any logic admitting reducible semantics”. This constructively preserves the essence of Completeness: provability in a geometric system exactly matches validity in all (classically adequate) semantic frameworks.
Definition 3.3.
Constructive arithmetic will be interpreted within a first-order predicative geometry $\mathcal{G}$ in the sense of Tarski [1954]. The language of $\mathcal{G}$ is governed by axioms of incidence, betweenness, and congruence:
$\text{(I1)}\;\forall x,y\,\exists z\,[B(x,z,y)\vee B(y,z,x)],$ (3.12)
$\text{(C1)}\;E(x,y,x,y),$ (3.13)
$\text{(C2)}\;E(x,y,u,v)\!\Rightarrow\!E(y,x,v,u),$ (3.14)
$\text{(T)}\;B(x,y,z)\wedge B(y,z,u)\!\Rightarrow\!B(x,y,u).$ (3.15)
Arithmetic predicates are represented by finite configurations: addition corresponds to segment concatenation, multiplication to rectangular composition, and equality to congruence.
Definition 3.4 (Geometric Semantics).
For terms $a,b,n\in\mathbb{N}$ let their geometric counterparts be finite sets $\Gamma(a)$, $\Gamma(b)$, $\Gamma(n)$ satisfying
$\operatorname{Rect}(a,b;n)\iff\exists P,Q,R,S\,\bigl[$ $B(P,Q,R)\wedge B(Q,R,S)\wedge{}$ (3.16)
$E(P,Q,R,S)\wedge\operatorname{A}(PQRS)=n\bigr].$
Thus $\mathsf{Comp}(n)$ asserts the existence of such a rectangle, and $\operatorname{Prime_{\Pi}}(n)$ its universal absence.
Exposition.
The realizability barrier corresponds both to the hardness of packability in geometric and to the absence of any uniform algorithm producing prime witnesses for all even integers. Under realizability semantics,
$\textsf{Provable in {HA}}\;\Longleftrightarrow\;\textsf{Realizable under BHK}\;\Longleftrightarrow\;\textsf{Representable in $\mathcal{G}.$}$
The model $\mathcal{G}$ denotes a constructive grid semantics in which bounded formulas $\phi(x)\in\Delta_{0}$ correspond to finite configurations $\Gamma_{\phi}(x)$, see [Tarski, 1954]. Existential ($\Sigma_{1}$) formulas describe the existence of such configurations; universal ($\Pi_{1}$) formulas assert that no configuration of the prescribed form exists. Arithmetic, logic, and geometry thereby appear as distinct projections of the same constructive semantics.
Definition 3.5 (Geometric Encoding).
For $n>1$ define
$\mathsf{Comp}(n)\iff\exists a,b>1\,\operatorname{Rect}(a,b;n),$ (3.17)
$\operatorname{Prime_{\Pi}}(n)\iff n>1\wedge\neg\mathsf{Comp}(n),$ (3.18)
where $\operatorname{Rect}(a,b;n)$ expresses the existence of a finite $a\times b$ tiling of total area $n$ within $\mathcal{G}$. The predicate $\operatorname{Rect}(a,b;n)$ is $\Delta_{0}$ and locally decidable; its existence provides a constructive witness of compositeness, see Figure 2. Conversely, $\operatorname{Prime_{\Pi}}(n)$ expresses the universal failure of such packings.
3.2 Failures of Bounded Local Definitions
Figure 3: The number 66 and 88 admit consistent rectangular decompositions, while 55 and 77 yield none. This illustrates the asymmetry of verification: composites produce explicit local witnesses (blue), whereas primality is detected only through the global ($\Pi_{1}$) absence of such configurations.
Observation.
Given an explicit rectangular packing, compositeness is decidable by a bounded constructive procedure. Let $\operatorname{Turn}_{3}(\gamma)$ denote a rectilinear loop $\gamma$ enclosing area $n$ with three right angles (the minimal closed rectangle). Define
$\Delta_{\square}(n)\;:\!\iff\;\exists a,b>1\,\exists\gamma\,[\operatorname{Rect}(a,b;n)\wedge\operatorname{Turn}_{3}(\gamma)].$ (3.19)
Then $\Delta_{\square}(n)$ is $\Delta_{0}$ and implies $\mathsf{Comp}(n)$. This traversal provides the geometric analogue of a $\Sigma_{1}$ witness: finite, decidable, and stable under substitution, see Figure 2. Every $n>1$ admits trivial strips $1\times n$ and $n\times 1$. Excluding these cases requires the universal constraint $a,b>1$, already a $\Pi_{1}$ clause. Hence determining that only trivial packings exist is nonlocal, see Figure 3 and Figure 5. For any finite bound $D$, there exist composites with all divisors >D>D, indistinguishable from primes under divisor bounds $\leq D$.
Figure 4: Primality as a logical Farey sequence problem, connecting to [Franel, 1924], on $xy=n$. Factor pairs within the finite bound $a,b\leq D$ (solid points) are detected by the bounded verifier, while those beyond the window remain invisible. Composites with large factors thus appear prime to any finite local schema—the Continuum Hypothesis of Cantor [1955] returns in a different form.
Theorem 3.6 (Collapse Patterns).
For every bounded local schema $d(a,b,n)$ defining compositeness, there exist “infinitely many” true composites not captured by it.
$\forall d\in\Delta_{0}\;\exists^{\infty}n\,[\,\mathsf{Comp}(n)\wedge\neg\mathsf{Comp}_{d}(n)\,].$ (3.20)
Proof Sketch.
By Diagonalization, see Subsection 5.3. ∎
Figure 5: Combinatorial growth in $n=9$. The sequence of structural configurations (gray) follows a fractal pattern until a composite witness appears (blue). As combinatorial complexity increases, local configurations become compressible.
Proof.
For any fixed divisor bound $D$, the formula $\mathsf{Comp}_{d}$ can only detect factors $a,b\leq D$. Composites with $a,b>D$ thus escape recognition, see Figure 6. Since the set of such composites is unbounded, no $\Delta_{0}$ schema achieves multiplicative completeness. ∎
Observation.
Enumerating slopes $a/b$ on the hyperbola $xy=n$ with denominators $b\leq D$ produces a “Farey-totality” through different aspect ratios, see Figure 4. Composites whose smallest factor exceeds $D$ remain undetected, see Figure 6, appearing prime to any bounded verifier. This provides the geometric analogue of the logical non-uniformization examined in Section 4. In the model $\mathcal{G}$, compositeness corresponds to the existence of a finite packing, and primality to the universal failure of such packings. No finite local rule bridges this asymmetry. The divisor predicate is therefore intrinsically non-local.
4 Core Logic
4.1 Packing Turing Machine as Constructive Realizer
Exposition.
We formalize the “packing” semantics, e.g., as in Figure 3, of $\mathsf{Comp}(n)$ by an explicit deterministic Turing Machine that halts on composite inputs and diverges on primes.
Logic 4.1.
Let
$M=(Q,\Sigma,\Gamma,\delta,q_{0},q_{\mathrm{acc}},q_{\mathrm{rej}})$ (4.21)
be a deterministic “three-tape” Turing Machine operating as follows:
  1. 1.
    Function.
    1. (a)
      Input: unary $1^{n}$ on Tape 1.
    2. (b)
      Output: $1^{a}\#1^{b}$ on Tape 3 if $n=ab$ with $a,b\geq 2$.
    3. (c)
      Behavior: $M$ halts iff $n$ is composite.
  2. 2.
    Structure. Tapes maintain the invariant forms
    $\text{Tape 1: }1^{n},\quad\text{Tape 2: }L\,A^{a}\,C\,a^{a}\,\mathdollar,\quad\text{Tape 3: }B^{b},$ (4.22)
    where $A$ encodes the current divisor candidate and $a$ its working copy. Each iteration tests divisibility by successive $a\geq 2$ via block-subtraction on Tape 1. When a full partition is found, one $B$ is appended per completed row, producing $n=ab$.
  3. 3.
    Transition scheme. The transition function $\delta$ is the finite union of routines:
    1. (a)
      Row count: consume $a$ symbols from Tape 1; if the block fits exactly, append one $B$ and reload the working zone.
    2. (b)
      Completion: if the scan exhausts $1^{n}$ with at least two rows, emit $1^{a}\#1^{b}$ and accept.
    3. (c)
      Failure: if a remainder remains, increment $a$ by one and restart.
    Undefined state pairs go to $q_{\mathrm{rej}}$ to keep $\delta$ total.
  4. 4.
    Correctness.
    • Soundness: Acceptance implies $n=ab$ with $a,b\geq 2$.
    • Completeness: For any composite $n$, some $a$ satisfies the halting condition.
    • $\Sigma_{1}$ behavior: For prime $n$, no $a$ divides $n$; the process diverges.
  5. 5.
    Complexity. For accepting inputs, $M$ performs $O(n)$ head moves and halts in linear time. Hence $M$ constitutes an explicit $\Sigma_{1}$ realizer for $\mathsf{Comp}(n)$, concretely implementing the geometric packing semantics in finite mechanical form.
4.2 Breakdown of Finite-Shape Schemata
Figure 6: Illustration of a bounded divisor lattice classifying primes (gold) while restricted to factors $\{2,3\}$. 25 (purple) is not recognized as composite by the local rule $\mathsf{Comp}_{2}(n)\!\equiv\!\exists a,b\!\in\!\{2,3\}\,[a\!\cdot\!b=n,a,b>1]$: The factors $(5,5)$ lie outside the enumerated lattice. This omission exemplifies how finite constructive systems lose multiplicative completeness once the factor bound is exceeded.
Definition 4.2 (Classical and Constructive Primality).
$\operatorname{Prime}(n)\iff(n>1)\wedge\forall a,b\!\leq\!n\,[a\!\cdot\!b=n\Rightarrow(a=1\vee b=1)],$ (4.23)
$\operatorname{Prime}$ is a $\Pi^{0}_{1}$ arithmetic predicate in $\mathsf{PA}$; $\operatorname{Prime_{\Pi}}$ is the corresponding $\Pi_{1}$ predicate under BHK semantics.
Definition 4.3 (Shape-Bounded Schema).
A finite geometric schema $d(a,b,n)$ encodes a bounded compositional rule:
$\mathsf{Comp}_{d}(n)\;\equiv\;\exists a,b\!\leq\!n\,[d(a,b,n)\wedge a,b>1\wedge a\!\cdot\!b=n]$ (4.24)
$\operatorname{Prime}_{d}(n)\;\equiv\;n>1\wedge\neg\mathsf{Comp}_{d}(n).$ (4.25)
The schema is governed by a bounding function $B(n)$ satisfying
$\forall n,a,b\!\leq\!n\,[d(a,b,n)\Rightarrow a\!\leq\!B(n)\ \vee\ b\!\leq\!B(n)].$ (4.26)
Lemma 4.4 (Finite Omission).
If $B(n)<\lfloor\sqrt{n}\rfloor$ for infinitely many $n$, then
$\forall d\,(\text{finite shape-bounded})\ \exists^{\infty}N\,[\mathsf{Comp}(N)\wedge\neg\mathsf{Comp}_{d}(N)].$ (4.27)
Proof Sketch.
A divisor a>B(n)a\!>\!B(n) with $a\!\cdot\!b\!=\!n$ escapes any bounded search below $B(n)$. Infinitely many $n$ possess such factors. If $B(n)$ eventually exceeds $\sqrt{n}$, the schema ceases to be finite: the verification cost of all potential $(a,b)$ pairs grows super-arithmetically. Thus no finite $d$ captures multiplicative completeness. See Friedman [1975, 1976]; Beeson [1985]. ∎
Exposition.
This quantitative limit mirrors a logical one: finite geometric omission corresponds to the failure of any uniform $\Sigma_{1}$ selector for $\Pi_{1}$ evidence. The geometric and logical barriers coincide under the Curry–Howard interpretation, see [Howard, 1980].
4.3 Reflection Principle
Proposition 4.5 (Non-Uniform Extraction).
By [Friedman, 1975; Beeson, 1985], let $\mathsf{R}(n,y)$ be a $\Sigma_{1}$ relation whose parameters include a $\Pi_{1}$ premise. No total primitive-recursive function $F$ satisfies
$\mathsf{HA}\vdash\forall n\,(\exists y\,\mathsf{R}(n,y)\Rightarrow\mathsf{R}(n,F(n))).$ (4.28)
Exposition.
Any such $F$ would be a total $\Sigma_{1}$ extractor for $\Pi_{1}$ evidence, contradicting normalization in $\mathsf{HA}$. Bounded geometric omission in Lemma 4.4 is the semantic image of this logical non-uniformity.
Lemma 4.6 (Diagonal Barrier).
By [Kleene, 1952; Howard, 1980], Assume a $\Sigma_{1}$ predicate $\operatorname{Dec}(e,n,m)$ decides $\operatorname{Prime_{\Pi}}(m)$. By the diagonal lemma, there exists $n_{e}$ such that
$\operatorname{Prime_{\Pi}}(n_{e})\leftrightarrow\neg\operatorname{Dec}(e,n_{e},n_{e}),$ (4.29)
implying,
$\mathsf{HA}\vdash\bot,$ (4.30)
Remark.
This diagonal argument parallels Rice [1953]: no total computable operator decides a nontrivial semantic property of partial functions. In $\mathsf{HA}$, $m$ is prime” plays that semantic role.
Proposition 4.7 (Reflection Principle).
For $\mathrm{GC}^{*}_{\Pi}$, either there exists a least counterexample or a uniform witness extractor $E$ exists. Since no such total $E$ can exist, the obstruction is logical:
$(\neg\mathrm{GC}^{*}_{\Pi})\Rightarrow\exists n_{0}\,(\neg\varphi(n_{0})\wedge\forall m<n_{0}\,\varphi(m)).$ (4.31)
Exposition.
This reflection isolates $\mathrm{GC}^{*}_{\Pi}$ on the boundary between constructive provability and arithmetical truth: the existence of a minimal counterexample is $\Sigma_{1}$, while uniform witness extraction is forbidden by the non-uniformization results above.
Definition 4.8 (Additive Convolution).
For $C\subseteq\mathbb{N}$ define
$(C\star C)(n)\;\equiv\;\exists p,q\!\leq\!n\,[C(p)\wedge C(q)\wedge p+q=n],$ (4.32)
a $\Sigma_{1}$ predicate describing additive closure.
Theorem 4.9 (Exactness).
If $C\subseteq\operatorname{Prime_{\Pi}}$ and
$\forall n,(\operatorname{Even}(n)\Rightarrow(C\star C)(n)),$ (4.33)
with $C$ $\Pi_{1}$-minimal among such sets, then $C(x)\Leftrightarrow\operatorname{Prime_{\Pi}}(x)$.
Proof Sketch.
Work in HA. The inclusion $C\subseteq\operatorname{Prime_{\Pi}}$ is given. For the converse, assume $\operatorname{Prime_{\Pi}}(p)$ and $p\notin C$. Consider the $\Pi_{1}$ set
$C^{\prime}\equiv{x:\ C(x)\ \vee\ x=p}.$ (4.34)
Since $x=p$ is $\Delta_{0}$ and $\operatorname{Prime_{\Pi}}(x)$ is $\Pi_{1}$, the predicate $C^{\prime}$ is $\Pi_{1}$, satisfies $C\subsetneq C^{\prime}\subseteq\operatorname{Prime_{\Pi}}$, and still has total even coverage because $C$ already has it. This contradicts the $\Pi_{1}$-minimality of $C$ unless $\operatorname{Prime_{\Pi}}\subseteq C$. Hence $C=\operatorname{Prime_{\Pi}}$. ∎
Remark.
The argument uses only: (i) closure of $\Pi_{1}$ under disjunction with $\Delta_{0}$ atoms, and (ii) that “total even coverage” is preserved when adding elements. Minimality forces equality.
Theorem 4.10 (Classical Translation).
By the Gödel–Gentzen Translation, the constructive non-realizability of $\mathrm{GC}^{*}_{\Pi}$ implies
$\mathsf{PA}\vdash\forall e,\exists n,\neg\Phi(e,n),$ (4.35)
where $\Phi(e,n)$ states that the $e$-th primitive-recursive operator realizes the $\Sigma_{1}$ witnesses of $\Pi_{1}$ primality required by $\mathrm{GC}^{*}_{\Pi}$.
Proof Sketch.
Assume, for contradiction, that
$\mathsf{PA}\vdash\exists e,\forall n,\Phi(e,n).$ (4.36)
By negative translation and soundness of the translation from PA to HA, we obtain
$\mathsf{HA}\vdash\neg\neg\exists e,\forall n,\Phi(e,n).$ (4.37)
“Markov-free” realizability converts this into a closed realizer term $t$ that, given $n$, produces $\Sigma_{1}$ witnesses demanded by $\Phi(e,n)$. By the Curry–Howard reading of $\mathrm{GC}^{*}_{\Pi}$, this $t$ is a uniform $\Sigma_{1}$ extractor for $\Pi_{1}$ primality, which contradicts non-uniformization in HA. ∎
Observation.
If Peano Arithmetic is consistent, it will deny existence of a total recursive index $e$ for our “Goldbach extractor”, while Heyting Arithmetic expresses the same limitation as a failure of normalization. The barrier translates across constructive and classical layers, but is interpreted differently. Bounded geometric schemata omit composites beyond $\sqrt{n}$; this quantitative defect induces non-uniformization in $\mathsf{HA}$. As usual, the diagonal lemma exposes the barrier syntactically.
5 Logical Induction
5.1 Inductive Inference and Semantic Reduction
Exposition.
We simulate classical logic within our constructive framework by inductive reasoning, examining structural correspondences. We appeal to the Theory of Inductive Inference, see Solomonoff [1964a, b] and to the Church–Turing Thesis: any realizible arithmetical predicate can be represented by a minimal constructive schema carrying equivalent realization. For primality,
$\operatorname{Prime_{\Pi}}(n)\;:\;\forall a,b\leq n\,[a\!\cdot\!b=n\Rightarrow(a=1\vee b=1)],$ (5.38)
the global quantifier alternation may be replaced by a local configuration rule on a finite constructive grid $\mathcal{G}$:
$\mathsf{Comp}(n)\;:\;\exists a,b>1\,[a\!\cdot\!b=n],\quad\operatorname{Prime_{\Pi}}(n)\;\Longleftrightarrow\;\neg\mathsf{Comp}(n).$ (5.39)
This correspondence preserves full logical content while eliminating extraneous ontological assumptions. We demonstrate two cases: First, primality can be characterized as a fragment of the Radó [1962] function, the Busy Beaver, over $\mathbb{N}$. Second, the Euler product can be recovered within fixed-point logic under the assumption that primes constitute a distributed diagonal structure.
5.2 Reduction to Beaver
Exposition.
Building on the “packing machine” $M$ of Subsection 4.1, whose halting encodes $\mathsf{Comp}(n)$, we now show that its operational semantics are “Turing-equivalent” to the classical halting predicate. We define a Prime Beaver as a “semantic reduction” of primality, thereby proving hardness internally rather than appealing to external complexity. The construction below should not be understood as a mere coding into a Busy Beaver framework.
Definition 5.1 (Prime Beaver).
We define the Prime Beaver as a Turing machine executing the transition system
[…]
Here $n$ is the number under test, $k$ the divisor candidate, and $S$ a finite “last in, first out” stack of pending subgoals. All operations are $\Delta_{0}$ decidable and primitive recursive in HA. The solving of one corresponds to trivially solving the other, and vice versa. The operational semantics of primes are emulating the behavior of the Busy Beaver function and coincide in degree. No encoding artifice is needed; the equivalence arises from the shared structure of $\Sigma_{1}$ termination and its constructive exhaustion. The machine halts with a nonempty stack $S$,
$S\iff\text{factorization found}\iff\text{(n) is composite.},$ (5.41)
hence,
$S=\emptyset\iff\text{$\nexists$ divisor}\geq\sqrt{(}n)\iff\text{$n$ is prime}.$ (5.42)
Definition 5.2 ($\mathrm{PB}$ Function).
Let $T(M)$ denote the runtime of a “packing machine” $M$. Define
$\mathrm{PB}(s)\;\triangleq\;\max\{\,T(M)\mid M\leq s\text{ states and halts}\,\}.$ (5.43)
Lemma 5.3.
Halting for $\mathrm{PB}$ machines is $\Sigma_{1}$ definable:
$\mathsf{Halts}_{\text{\rm PM}}(M)\;\equiv\;\exists t\,\Theta(M,t),$ (5.44)
where $\Theta$ is a bounded arithmetical predicate expressing that $M$ reaches a terminal configuration after $t$ transitions.
Proof Sketch.
Unroll the computation as a finite sequence of configurations
[…]
Verification of each step is $\Delta_{0}$, so the overall existence claim is $\Sigma_{1}$. ∎
Lemma 5.4 (Upper Bound).
There exists a primitive-recursive compiler $\mathsf{Comp}$ and constant $c$ such that
$T(\mathsf{Comp}(M))\;\geq\;T(M),\quad\#\text{states}(\mathsf{Comp}(M))\leq c\cdot\#\text{states}(M).$ (5.46)
Hence
$\mathrm{PB}(s)\;\leq\;\mathrm{BB}(c\cdot s).$ (5.47)
Lemma 5.5 (Lower Bound).
There exist primitive-recursive functions $r,r^{\prime}$ and an injective map $M\mapsto P_{M}$ from Turing machines to “prime number machine” such that
$\#\text{states}(P_{M})\leq r^{\prime}(\#\text{states}(M)),$ (5.48)
$\neg\mathsf{Halts}(M)\Rightarrow T(P_{M})\leq C,$
[…]
where $F$ is a strictly increasing primitive-recursive padding function exceeding any total recursive bound prescribed in advance.
Theorem 5.6 (Semantic Reduction).
Combining Lemmas 5.4 and Lemma 5.5,
$\mathrm{PB}\equiv_{T}\mathrm{BB},\quad\deg_{T}(\mathrm{PB})=0^{\prime},$ (5.49)
and
$\forall f\in\mathrm{Rec}\;\exists s_{0}\;\forall s\geq s_{0}:\;\mathrm{PB}(s)>f(s).$ (5.50)
Thus $\mathrm{PB}$ dominates every total recursive function and attains the same Turing Degree as the classical Busy Beaver function.
5.3 Divisor Impredicativity
Heuristic 5.7 (Primes as Natural Diagonals).
Enumerate all rational aspect ratios corresponding to finite rectangular packings,
$\mathcal{R}=\left\{\frac{1}{1}\;,\;\frac{1}{2}\;,\;\frac{1}{3}\;,\;\dots\right\},$ (5.51)
each ratio encoding a possible composite decomposition ($n=a\times b$) through ($a/b$). The enumeration defines a countable grid of candidate factorizations. Constructively, every listed ratio corresponds to a realized composite; however, by Cantor’s Diagonal Argument [Cantor, 1891], one may define a new irrational slope not represented in any row or column of the enumeration. This “diagonal slope” corresponds to an integer configuration that eludes all finitely coded factor pairs—precisely the constructive analogue of a prime. Thus, the diagonal argument reproduces within arithmetic the same structural phenomenon as in the Continuum Hypothesis: the totality of countable constructive packings fails to exhaust the uncountable space of potential divisive configurations. Primality marks the diagonal complement of enumerated composites, the logical remainder that cannot be captured by any single recursive list of ratios.
Heuristic 5.8 (Cohen Forcing Analogue for Primality).
Let $\mathcal{M}\vDash\mathsf{HA}$ be a constructive model of arithmetic, and define forcing conditions as finite partial factorizations:
$p={(a_{i},b_{i}):a_{i}b_{i}=n_{i},;a_{i},b_{i}>1},$ (5.52)
ordered by extension $q\leq p$ if $q$ extends every pair in $p$. A “generic filter” $G\subseteq\mathbb{P}$ represents a consistent infinite family of such decompositions. The forcing language expands $\mathcal{L}_{A}$ by a predicate $\dot{F}$ meaning “is in the generic factorization.” Interpretation proceeds by recursion:
$p\Vdash\mathsf{Comp}(\bar{n})$ $\iff\exists(a,b)\in p,\quad[ab=n],$ (5.53)
$p\Vdash\operatorname{Prime_{\Pi}}(\bar{n})$ $\iff\forall q\geq p\neg\exists(a,b)\in q,\quad[ab=n].$
“Forcing” thus constructs an “arithmetical extension” preserving all $\mathsf{HA}$ theorems but deciding additional composite instances—exactly as the original independence proof of Cohen [1963, 1964], preserved $\mathsf{ZFC}$ while deciding $\neg\mathsf{CH}$.
Theorem 5.9 (Cohen Extension for Arithmetic).
If $\mathsf{HA}$ is consistent, there exist generic extensions $\mathcal{M}[G_{1}]$ and $\mathcal{M}[G_{2}]$ such that
$\mathcal{M}[G_{1}]\vDash\mathrm{GC}^{*}_{\Pi},\quad\mathcal{M}[G_{2}]\vDash\neg\mathrm{GC}^{*}_{\Pi}.$ (5.54)
Exposition.
Cohen’s argument generalizes: Forcing in set-theory adds new subsets of $\omega$; Forcing here adds new $\Sigma_{1}$ witnesses (partial decompositions). Both preserve consistency while extending definability. In this light, primality corresponds to a “generic gap”—a number immune to all extensions by finite factorizations—mirroring how new reals are adjoined beyond any countable ground model in $\mathsf{ZFC}$.
5.4 Euler Product from Fixed-Point
Exposition.
If the property of primality depends solely on an ontological postulate, it can be reproduced (incompletely) without such an assumption by syntactic simulation within a constructive formal system. Conversely, if primality manifests intrinsic impredicativity—its definition referring implicitly to the totality of potential factors—then it can be encoded (completely) as a distributed fixed point of constructive operators by Bekić’s Lemma [Jones and Bekić, 1984]. Within this formulation, the Euler Product arises not as an analytic artifact but as the diagonal solution of a purely logical recursion, derived in fixed-point logic without arithmetic premises.
Definition 5.10 (Bekić’s Lemma).
Let $(D,\leq)$ be a complete lattice and $f,g:D\times D\to D$ be monotone functions. Define
$F(x,y)=(f(x,y),g(x,y)).$ (5.55)
If the least fixed point $\mu y.\,g(x,y)$ exists for each $x$, then the least fixed point of $F$ satisfies
$\mu(x,y).\,F(x,y)=\bigl(\mu x.\,f(x,\,\mu y.\,g(x,y)),\;\mu y.\,g(\mu x.\,f(x,\,\mu y.\,g(x,y)),y)\bigr).$ (5.56)
Bekić’s Lemma thus decomposes joint least fixed points into nested single-variable ones.
Logic 5.11 (Simulation of the Euler Product).
Assume within HA that primality is a $\Pi_{1}$ predicate lacking any total $\Sigma_{1}$ extractor, i.e.,
$\neg\exists E\,\forall n\,[\operatorname{Prime_{\Pi}}(n)\Rightarrow\exists a\leq E(n)\,R(n,a)].$ (5.57)
This non-realizability establishes a fixed-point constraint on the constructive operators $R_{\Sigma}$ and $R_{\Pi}$:
$R_{\Sigma}R_{\Pi}(x)=x\quad\textsf{has no constructive solution in\penalty 10000\ $\mathsf{HA}$.}$ (5.58)
Under Bekić’s Lemma, the least fixed point of the compositeness operator
$\mathsf{Comp}(n)=\mu_{X}\,\exists a,b\,[a,b>1\wedge(a\!\cdot\!b=n\vee X(a)\vee X(b))]$ (5.59)
yields a canonical separation of the multiplicative structure into independent prime components. Summation of these decomposed layers produces the definitional identity
$\zeta(s)=\sum_{n\geq 1}n^{-s}=\prod_{p\in\mathcal{P}}\frac{1}{1-p^{-s}}.$ (5.60)
Proposition 5.12.
Within HA, let $R_{\Sigma}$ and $R_{\Pi}$ denote constructive closure under bounded existential and universal search over $\Delta_{0}$ formulas. For a base predicate $\varphi(x)\in\Delta_{0}$, define
$R_{\Sigma}(\varphi(x))\equiv\exists y\!\leq\!x\,\varphi(y),\quad R_{\Pi}(\varphi(x))\equiv\forall y\!\leq\!x\,\varphi(y).$ (5.61)
Iterating these operators alternately yields a sequence of definitional layers:
$\varphi_{n+1}(x)\equiv R_{\Sigma}(R_{\Pi}(\varphi_{n}(x))).$ (5.62)
Each iteration introduces one alternation of constructive witnessing and reflective verification. Let the meta-analytic ratio of successive definitional complexities be denoted by
$\frac{\varphi_{n+2}}{\varphi_{n+1}}=1+\frac{1}{\frac{\varphi_{n+1}}{\varphi_{n}}}.$ (5.63)
Its fixed point is the constant $\Phi$, which quantifies the equilibrium between constructive expansion ($\Sigma_{1}$ search) and reflective contraction ($\Pi_{1}$ verification). See Section Context for related discussion.
Thesis 5.13 ($\Phi$-Adic Equilibrium).
Consider the fixed-point schema
$R_{\Sigma}R_{\Pi}(x)=x.$ (5.64)
HA cannot prove the existence of a solution:
$\textsf{HA}\nvdash\exists x\,[R_{\Sigma}R_{\Pi}(x)=x].$ (5.65)
This expresses the failure of constructive evidence under normalization. Realization of such a fixed point would require a non-constructive operator; the system thereby enforces its own limit of definability.
Proposition 5.14 (Recursive Depth Law).
Let $n_{k}$ denote successive bounds arising from alternating $\Sigma_{1}/\Pi_{1}$ elimination on a bounded arithmetic predicate. If $n_{k+1}=\lfloor\sqrt{n_{k}}\rfloor$ with $n_{0}=n$, termination occurs after
$k(n)\sim\log_{\!\Phi}n.$ (5.66)
Remark.
The constant $\Phi$ quantifies the contraction rate between local constructive computation and global reflective closure. It “measures” the ratio at which definitional alternation stabilizes: existential search expands complexity, while universal verification contracts it. Their alternation converges toward the slowest transformation [Hurwitz, 1891], marking the constructive boundary where witnessing and verification reach equilibrium—“after the next step.”
5.5 Decomposition and Factorization
Proposition 5.15 (Prime Generation).
Let $(\mathbb{N},\cdot)$ be the free commutative monoid on the prime generators $\mathcal{P}$. For $s\in\mathbb{C}$ with $\Re(s)>1$, define the multiplicative weight
$\mu_{s}(n)\equiv n^{-s}.$ (5.67)
Dirichlet addition is ordinary summation, and multiplication acts additively on exponents, see [Apostol, 1976]:
$\mu_{s}(ab)=\mu_{s}(a)\mu_{s}(b).$ (5.68)
Definition 5.16 (Constructive fixed-point Formulation).
Define the compositeness operator
$\mathsf{Comp}(n)=\mu_{X}\,\exists a,b\,[a,b>1\wedge(a\!\cdot\!b=n\vee X(a)\vee X(b))],$ (5.69)
whose least fixed-point $\mu_{X}$ captures closure under multiplicative decomposition. By Bekić’s Lemma, the joint recursion decomposes coordinatewise:
$\mu_{(a,b)}F(a,b,n)=\big\langle\mu_{a}(\exists b\,F(a,b,n)),\;\mu_{b}(\exists a\,F(a,b,n))\big\rangle.$ (5.70)
Each coordinate corresponds to a prime component $p\in\mathcal{P}$, yielding the constructive prime lattice.
Proposition 5.17 (Prime-Power Slices).
For each prime $p\in\mathcal{P}$ define
$\mathcal{M}_{p}=\{p^{k}\mid k\geq 0\},\quad\mu_{s}|_{\mathcal{M}_{p}}(p^{k})=p^{-ks}.$ (5.71)
Every $n\in\mathbb{N}$ corresponds uniquely to a finitely supported function
$\nu_{p}:\mathcal{P}\!\to\!\mathbb{N}$ (5.72)
with
$n=\prod_{p}p^{\nu_{p}},$ (5.73)
so that
$\mathbb{N}\;\simeq\;\bigsqcup_{(\nu_{p})\ \text{finite}}\;\prod_{p\in\mathcal{P}}\mathcal{M}_{p}.$ (5.74)
Theorem 5.18 (Constructive Factorization of $\zeta$).
Let $\Re(s)>1$ and define
$\zeta(s)\equiv\sum_{n\geq 1}\mu_{s}(n).$ (5.75)
Then by Bekić’s Lemma and decomposition on product lattices,
$\zeta(s)=\sum_{(\nu_{p})\ \text{finite}}\ \prod_{p\in\mathcal{P}}p^{-\nu_{p}s}=\prod_{p\in\mathcal{P}}\Big(\sum_{k\geq 0}p^{-ks}\Big)=\prod_{p\in\mathcal{P}}\frac{1}{1-p^{-s}}.$ (5.76)
Proof Sketch.
Since $\mu_{s}$ is completely multiplicative,
$\mu_{s}\!\Big(\prod p^{\nu_{p}}\Big)=\prod p^{-\nu_{p}s}.$ (5.77)
Summing over all finitely supported exponent vectors $(\nu_{p})$ yields the product of geometric series
$\sum_{k\geq 0}p^{-ks}.$ (5.78)
Absolute convergence for $\Re(s)>1$ allows interchange of product and sum, producing the classical Euler Product. ∎
Corollary 5.19 (Logarithmic Derivative).
For $\Re(s)>1$,
$\frac{\zeta^{\prime}(s)}{\zeta(s)}=\sum_{p\in\mathcal{P}}\frac{\log p}{p^{s}-1}=\sum_{n\geq 1}\frac{\Lambda(n)}{n^{s}},$ (5.79)
where $\Lambda$ denotes the Von Mangoldt Function, see [von Mangoldt, 1895].
Remark.
The factorization is constructive up to the analytic step of convergence. Within HA, the bisection already entails unique multiplicative structure.
Observation.
The Euler Product defines $\zeta(s)$ for $\Re(s)>1$. Classically, $\zeta(s)$ extends meromorphically to $\mathbb{C}$ and satisfies
$\zeta(s)=2^{s}\pi^{s-1}\sin\!\left(\frac{\pi s}{2}\right)\Gamma(1-s)\zeta(1-s).$ (5.80)
The reflection $s\mapsto 1-s$ acts as the analytic analogue of the $\Sigma_{1}/\Pi_{1}$ duality. This transformation re-expresses the unprovable fixed-point schema
$R_{\Sigma}R_{\Pi}(x)=x$ (5.81)
In this sense,
$\mathsf{HA}\quad\Longleftrightarrow\quad\textsf{Continuum simulation}\quad\Longleftrightarrow\quad\textsf{Zeta approximation.}$
The fixed-point reconstruction therefore aligns the logical boundary of constructive definability of $\zeta(s)$. Whether prime distribution reflects numerical contingency or logical necessity depends on how one interprets this boundary between constructive structure, arithmetical truth and simulation.
6 Corollaries and Implications
6.1 Witnessing Form
Definition 6.1 (Goldbach Witness).
Let $\operatorname{Cert}(p,w)$ be a $\Sigma_{1}$ predicate satisfying
$\mathsf{PA}\vdash\forall p\,\big(\operatorname{Prime_{\Pi}}(p)\leftrightarrow\exists w\,\operatorname{Cert}(p,w)\big).$ (6.82)
Define the witnessing form
$\mathrm{WGC}:\;\forall n\Big(\operatorname{Even}(n)\Rightarrow\exists p,q,w_{1},w_{2}\,[\operatorname{Cert}(p,w_{1})\wedge\operatorname{Cert}(q,w_{2})\wedge p+q=n]\Big).$ (6.83)
Lemma 6.2 (Classical Equivalence).
$\mathsf{PA}\vdash\mathrm{GC}^{*}_{\Pi}\;\leftrightarrow\;\mathrm{WGC}.$ (6.84)
Proof Sketch.
Replace each occurrence of $\operatorname{Prime_{\Pi}}(p)$ in $\mathrm{GC}^{*}_{\Pi}$ by $\exists w\,\operatorname{Cert}(p,w)$ and simplify. The equivalence follows from propositional reduction inside $\mathsf{PA}$. ∎
Exposition.
Classically, $\mathrm{GC}^{*}_{\Pi}$ and $\mathrm{WGC}$ coincide since $\Sigma_{1}$ witnessing introduces no proof cost. Constructively, however, $\mathrm{WGC}$ presupposes a total recursive Skolem function producing $(p,q,w_{1},w_{2})$ uniformly— the very operation prohibited by non-uniformization in $\mathsf{HA}$ [Friedman, 1976; Beeson, 1985; Troelstra and van Dalen, 1988].
6.2 Stratification
Logic 6.3 (Provability Stratification).
We conduct case analysis:
$\begin{array}[]{rll}(i)&\mathsf{PA}\vdash\mathrm{GC}^{*}_{\Pi},&\text{true and provable},\\[2.0pt] (ii)&\mathsf{PA}\vdash\neg\mathrm{GC}^{*}_{\Pi},&\text{false and refutable},\\[2.0pt] (iii)&\mathsf{PA}\nvdash\mathrm{GC}^{*}_{\Pi},\ \mathsf{PA}\nvdash\neg\mathrm{GC}^{*}_{\Pi},&\text{independent.}\end{array}$ (6.85)
Constructive analysis eliminates case (i): even if $\mathrm{GC}^{*}_{\Pi}$ holds in $\mathbb{N}$, it lacks a total $\Sigma_{1}$ constructor in $\mathsf{HA}$. Hence:
$\mathrm{Thm}_{\mathsf{HA}}\subsetneq\mathrm{Thm}_{\mathsf{PA}}\subsetneq\mathrm{True}_{\mathbb{N}}.$
Remark.
Finite geometric schemata capture only bounded multiplicativity. Thus $\mathsf{HA}$ fails to uniformize $\Pi_{1}$ predicates, and $\mathsf{PA}$ reflects this as the nonexistence of a total recursive witness index. The constructive barrier delineates the strength gap between $\mathsf{HA}$ and $\mathsf{PA}$.
6.3 Criterion
When the $\mathsf{HA}$ obstruction arises from a uniform, total realizer of a $\Pi_{1}\!\Rightarrow\!\Sigma_{1}$ relation, it transfers to $\mathsf{PA}$ as “no total recursive index exists”.
Corollary 6.4.
Let $R(n,y)$ be $\Sigma_{1}$. For a (partial) recursive function $\varphi_{e}$, write
$\Phi(e,n)\;:\;\big(\Pi_{1}(n)\Rightarrow\exists y\,R(n,y)\big)\;\wedge\;\big(\varphi_{e}(n)\!\downarrow\Rightarrow R\big(n,\varphi_{e}(n)\big)\big).$ (6.86)
If, in $\mathsf{HA}$, there is no total operator extracting witnesses,
$\mathsf{HA}\vdash\neg\exists F\,\forall n\big(\Pi_{1}(n)\Rightarrow R(n,F(n))\big),$ (6.87)
then
$\mathsf{PA}\vdash\forall e,\exists n\,\neg\Phi(e,n).$ (6.88)
Equivalently, $\mathsf{PA}$ proves that no recursive index $e$ works uniformly.
Exposition.
The constructive results established in Section 4 delimit realizability within $\mathsf{HA}$. We now trace the classical reflections of this barrier in $\mathsf{PA}$, where normalization no longer constrains proofs, but recursion-theoretic and proof length limits persist. A prime conjecture admits constructive resolution within $\mathsf{HA}$ iff its arithmetical form allows total witness extraction by a primitive-recursive operator. Formally, for
$\forall n\,\big(\Pi_{1}(n)\Rightarrow\exists y\,\Sigma_{1}(n,y)\big),$ (6.89)
constructive realizability requires a computable bound $B(n)$ such that all admissible witnesses $y$ satisfy $y\!\leq\!B(n)$ independently of the $\Pi_{1}$ premise. If such a bound exists, the operator
$E(n)\;\equiv\;\mu y\!\leq\!B(n)\,[\Sigma_{1}(n,y)]$ (6.90)
is total and provably terminating in $\mathsf{HA}$. If no such bound exists, any total extractor $E$ would decide a nontrivial $\Pi_{1}$ property, violating normalization. Hence,
$\textsf{Witnesses}\;\Rightarrow\;\textsf{Realization},\quad\textsf{$\Pi_{1}$ dependence}\;\Rightarrow\;\textsf{Obstruction}.$
Theorem 6.5 (Classification).
Let $\mathcal{C}$ denote an arithmetical conjecture expressible in $\mathsf{PA}$.
  1. (i)
    If $\mathcal{C}$ has a decidable witness domain with a primitive-recursive bound, then $\mathsf{HA}\vdash\mathcal{C}$ (constructively provable).
  2. (ii)
    If $\mathcal{C}$ involves nonuniform $\Pi_{1}$ dependence without bounded search, then $\mathsf{HA}\nvdash\mathcal{C}$. However, once $\mathcal{C}$ is classically reformulated by invoking the Law of the Excluded Middle or double-negation elimination, it may become provable in $\mathsf{PA}$:
    $\mathsf{PA}\vdash\mathcal{C}\quad\text{while}\quad\mathsf{HA}\nvdash\mathcal{C}.$ (6.91)
    This translation replaces constructive witnessing by classical truth valuation, converting a nonuniform $\Pi_{1}$ dependence into a closed arithmetic predicate admissible classically.
Example 6.6 (Fermat’s Equation).
$\nexists n(x^{n}+y^{n}=z^{n},n>2,n\in\mathbb{N}),$ (6.92)
defines a $\Pi_{0}$ predicate: its falsity is verified by a finite search over bounded integers once an explicit counterexample domain is fixed. The constructive proof of its negation requires only the absence of such bounded witnesses, a $\Sigma_{1}$ refutation compatible with normalization in $\mathsf{HA}$. By contrast, the constructive Goldbach principle
$\mathrm{GC}^{*}_{\Pi}:\;\forall n\big(\operatorname{Even}(n)\Rightarrow\exists p,q\,[\operatorname{Prime_{\Pi}}(p)\wedge\operatorname{Prime_{\Pi}}(q)\wedge p+q=n]\big)$ (6.93)
involves unbounded $\Pi_{1}$ dependence: the primality clauses require verification over all possible divisors. Any uniform extraction of $(p,q)$ would thus decide a nontrivial $\Pi_{1}$ predicate, violating normalization in $\mathsf{HA}$. Hence Fermat’s Theorem remains constructively decidable, while Goldbach’s Conjecture constructive form exceeds the realizable limits of $\mathsf{HA}$.
Example 6.7 (Helfgott’s Proof).
The ternary Goldbach conjecture (every odd $n>5$ is a sum of three primes) is provable in classically, following Helfgott [2013]. Its arithmetical form is $\Pi^{0}_{2}$ with a decidable witness domain, hence lies within case (i) of the stratification. It introduces no uniform $\Pi_{1}$ extraction and therefore does not violate the constructive barrier.
6.4 Oracle Violation
Definition 6.8 (Uniform Skolem Function).
Assume a total primitive–recursive operator $E$ satisfying
$\operatorname{Even}(\bar{n})\Rightarrow\bigl[$ $\operatorname{Cert}(\pi_{1}E(\bar{n}),\pi_{3}E(\bar{n}))\wedge{}$ (6.94)
$\operatorname{Cert}(\pi_{2}E(\bar{n}),\pi_{4}E(\bar{n}))\wedge{}$
$\pi_{1}E(\bar{n})+\pi_{2}E(\bar{n})=\bar{n}\bigr].$
Lemma 6.9 (Definitional Extension).
Extending $\mathsf{PA}$ by the operator $E$ yields a definitional theory $\mathsf{PA}(E)$. For each numeral $\bar{n}$, the instance of (6.94) is derivable by substitution within $\mathsf{PA}(E)$.
Corollary 6.10 (Uniform Proof Length).
If (6.94) held, then each $\mathrm{WGC}(\bar{n})$ would admit a proof in $\mathsf{PA}(E)$ of constant length independent of $n$.
Theorem 6.11 (Oracle Violation).
No such operator $E$ exists.
Proof Sketch.
By the Gödel–Blum speedup theorems [Gödel, 1986; Blum, 1967], no arithmetical theory extending $\mathsf{PA}$ can have uniformly bounded proofs for an infinite family of true sentences. A total $E$ would induce such uniformity, collapsing the proof-length hierarchy. Hence $E$ cannot exist. ∎
Observation.
A uniform extractor $E$ would act as a logical oracle, erasing the boundary between instance and derivation. The speedup theorem forbids this compression. Constructively, the prohibition manifests as the nonexistence of a total $\Sigma_{1}$ extractor; classically, as the absence of a Skolem function collapsing proofs.
7 Interpretations
This section offers a speculative interpretation of the constructive and logical results obtained above through a semantic reduction linking realizability, computation, and informational structure.
7.1 Excluded Middle
Heuristic 7.1 (Divisor Instances as LEM).
No constructive perspective is complete without discussing the law of the excluded middle, or LEM.
$\textsc{LEM}_{\Gamma}:\quad\forall P\in\Gamma\,[P\vee\neg P].$ (7.95)
HA proves $\textsc{LEM}_{\Delta_{0}}$ but not $\textsc{LEM}_{\Sigma_{1}}$ or $\mathrm{LEM}_{\Pi_{1}}$. The predicates
$\mathsf{Comp}(n):\exists a,b>1\,[a\!\cdot\!b=n],$ (7.96)
$\operatorname{Prime_{\Pi}}(n):\forall a,b\leq n\,[a\!\cdot\!b=n\Rightarrow(a=1\vee b=1)].$ (7.97)
illustrate this asymmetry. $\mathsf{Comp}(n)$ is $\Sigma_{1}$ and finitely decidable; $\operatorname{Prime_{\Pi}}(n)$ is $\Pi_{1}$ and constructively undecidable. Without $\mathrm{LEM}_{\Pi_{1}}$, HA cannot prove $\mathsf{Comp}(n)\vee\operatorname{Prime_{\Pi}}(n)$ for arbitrary $n$. Any uniform constructor transforming $\Pi_{1}$ primality into $\Sigma_{1}$ witnesses would constitute a total realization of $\mathrm{LEM}_{\Pi_{1}}$, contradicting normalization in HA.
7.2 Primes as Proof of Knowledge
Thesis 7.2 (Rigged System).
If a thermodynamic, computational, or informational model contains internally conserved or limitative quantities, then their distribution expresses a constraint of relative consistency rather than a contingent pattern accessible to finite derivation. The quantities conventionally termed “energy”, “information”, or “entropy” correspond to proofs measured against descriptive cost. Any global regularity governing them thus represents a conservation law at the “meta-level”; its proof would require transcending the system’s closure, thereby violating the very conservation it asserts.
Observation.
For example, Quantum computation accelerates instance evaluation but does not transform a given nonconstructive predicate into a realization. Logical non-uniformity persists across such computation, see [Abramsky, 2004].
Heuristic 7.3 (Constructive Reflection in Quantum Algorithms).
For modular exponentiation $f(x)=a^{x}\bmod N$, determining the order $r$ such that
$a^{r}\equiv 1\pmod{N}\quad\text{and}\quad\forall k<r\,[a^{k}\not\equiv 1\pmod{N}$]
is a mixed $\Pi_{1}$-$\Sigma_{1}$ task. Shor’s algorithm [Shor, 1994; Ekert and Jozsa, 1996] achieves an instancewise realization by constructing a unitary map
$U_{f}:|x\rangle\mapsto|f(x)\rangle$ (7.98)
and applying the quantum Fourier transform $\mathcal{F}$. Measurement yields one valid witness $r$ per run, but no total operator $E:\mathbb{N}\!\to\!\mathbb{N}$ valid for all $N$.
Heuristic 7.4 (Entropy–Packing Correspondence).
Entropy can be understood as a constructive packing problem: The computational hardness of identifying such minimal states and the thermodynamic cost of reducing entropy are therefore relatively consistent in respect to a shared structural constraint. Let $W_{n}$ denote the number of admissible $\Delta_{0}$ tilings (finite constructive configurations) of area $n$. Define the combinatorial entropy
$S_{n}=k\log W_{n},$ (7.99)
in Boltzmann’s sense, where $k$ is a normalizing constant. If successive configuration counts satisfy
$\frac{W_{n+2}}{W_{n+1}}=1+\frac{1}{\frac{W_{n+1}}{W_{n}}},$ (7.100)
then
$\frac{S_{n+2}-S_{n+1}}{S_{n+1}-S_{n}}\longrightarrow\Phi,$ (7.101)
expressing $\Phi$-proportional growth and minimal informational curvature in the sense of Jaynes [1957]; Landauer [1961].
Exposition.
(Analogy.) Consider a finite room scattered with boxes of various sizes. A computer is tasked with determining an optimal packing configuration. As the algorithm explores permutations, we observe from the outside that each increment in disorder—i.e., each deviation from a constructive arrangement—raises its computational energy cost. Entropy increases with unordering, but so does the cost of realization. Now imagine embedding this simulation within another computer that, in turn, simulates the first. The inner system must entail the minimal cost of fluctuations by necessity, accumulating an additional layer of informational entropy.
Heuristic 7.5 (Harmonic Log and Energy Gradient).
Let the potential
$H(x)=\sum_{p\leq x}\frac{1}{p}$ (7.102)
measure the aggregate informational weight of primes up to $x$, diverging as
$H(x)\sim\log\log x.$ (7.103)
Define the local prime–composite ratio
$\rho(x)\;=\;\frac{\pi(x)}{x-\pi(x)}\;\approx\;\frac{1}{\ln x-1},$ (7.104)
which acts as an occupation fraction: the probability that a natural number occupies a “prime state”. The gradient
$\frac{d\rho}{dx}\;\approx\;-\frac{1}{x(\ln x-1)^{2}}$ (7.105)
then expresses the decay of that occupation with energy $x$, directly analogous to the Boltzmann Transport Equation
$\frac{dn(E)}{dE}\propto-\frac{n(E)}{kT},$ (7.106)
where $n(E)$ is the particle density over energy $E$. Thus, the prime distribution behaves as an entropy-limited flow: primes correspond to “knowledge” (in a limitative sense), since each prime implicitly enumerates both itself and the composites it generates, while composites merely instantiate that prior structure. Hence, knowing composites conveys only local configurations, but knowing primes entails global knowledge of both classes—an epistemically costlier act.
Heuristic 7.6 (Cost of Knowledge).
In any formal system $\mathcal{F}$ that encodes Turing-complete arithmetic and remains relatively consistent with a stronger system $\mathcal{F}^{\prime}$ conserving an invariant $Q$, the realization of a new prime $p$ (i.e. a new irreducible witness) requires an increase in descriptive or computational cost proportional to its informational deficit:
$\mathrm{Cost}_{\mathcal{F}}(p)\;\propto\;-\frac{\partial S_{n}}{\partial n}\Big|_{n=p}\;\approx\;\frac{k}{\ln p}.$ (7.107)
Hence primes emerge as proofs of knowledge: each encodes compression relative to its compositional background, and their discovery or verification is computationally expensive within any consistent finitary system.
Remark.
The correspondence is empirically motivated but structural: Once we axiomatize that inference incurs energetic cost, then—without assuming informational substrate—there exist a conserved quantity. In any arithmetical or computational framework subject to a conserved measure of informational entropy, the (non-probabilistic) generation of new prime predicates corresponds to maximal local decrease in entropy and thus maximal proof cost. Primes mark points where the system attains minimal redundancy.
8 Discussion
We conclude with the operational semantics of computation. Primality in computer science is implemented (usually) by bounded search:
$\mathsf{Prime}_{d}(n)\;:\;\forall a\leq d(n)\,[a\mid n\Rightarrow(a=1\vee a=n)],$ (8.108)
with $d(n)$ a fixed bound such as $\lfloor\sqrt{n}\rfloor$ or a bit-length constraint. Such predicates are $\Delta_{0}$ because all quantifiers range over explicit finite domains. Once division is realized by iterative subtraction, modular approximation, or simulated quotient extraction, it becomes a surrogate for a non-constructive relation [Cutland, 1980]. Hence every modern system interprets $\operatorname{Prime_{\Pi}}(n)$ as a bounded, effectively decidable predicate. Testing thereby becomes an evaluation of a finite schema, not a $\Pi_{1}$ reflection.
Let $\mathcal{M}_{e}$ denote the $e$-th Turing machine. For each $\mathcal{M}_{e}$, there exists $p_{e}$ such that $\mathcal{M}_{e}$ neither halts affirming nor refuting $\operatorname{Prime_{\Pi}}(p_{e})$:
$\forall e\,\exists p_{e}\;\neg\big(\mathcal{M}_{e}(p_{e})\!\downarrow\ \wedge\ (\mathcal{M}_{e}(p_{e})=\text{p}\ \vee\ \mathcal{M}_{e}(p_{e})=\text{c})\big),$ (8.109)
a direct consequence of the non-uniformization theorem [Friedman, 1975] and Rice’s diagonal argument [Rice, 1953].
This computational constraint mirrors the constructive one. The logical asymmetry analyzed throughout—from local tilings to global indivisibility—reappears as a fundamental limit of mechanized arithmetic. Constructive logic, proof theory and computation share a single invariant: there always exists a prime beyond any machine’s operational horizon, owing to the fact that modern computation emulates division by bounded iteration of its subsystems; it never realizes the relation $\operatorname{div}(a,b)$ in total form.
$\exists(a,b)\;\neg\big(\mathcal{M}_{e}(a,b)\!\downarrow=\operatorname{div}(a,b)\big).$ (8.110)
Thus every algorithmic system remains a partial interpreter of arithmetic, not its total executor. Many classical theorems—e.g., Euclid’s Theorem and termination statements for recursive systems—share the form
$\forall x\,\exists y\,R(x,y).$ (8.111)
Constructively, such claims rely on unbounded meta-quantification. Their truth expresses an appeal to infinity: a limit process that no finite computation can exhaust. Proving a $\Pi_{2}$ statement does not close the infinite but formalizes its containment within arithmetic:
Each quantifier alternation $\forall x\,\exists y$ achieves compression
through the semantics of its own incompleteness.
No $\Pi_{2}$ problem has ever been resolved by purely finite means without invoking a transfinite assertion. This transition is not a failure but a trade: resolving obstructive problems under subsumation that preserves consistency under uncertainty entails an approximation of
$\forall x\,\exists y\,R(x,y),$ (8.112)
that again admits compression by bounded schemas,
$\forall x\leq D\,\exists y\leq B(x)\,R(x,y),$ (8.113)
where $D$ and $B(x)$ mark the constructive horizon of verification.
Our result locates the hardness of any prospective proof of Goldbach’s Conjecture—and, by extension, the constructive status of primality—within the finitist and predicativist tradition. The obstacle is not primarily analytic or combinatorial but logical: it arises at the boundary of constructive definability, where any attempt to furnish a uniform $\Sigma_{1}$ extractor for $\Pi_{1}$ predicates encounters the same limitations described in ultra-finitistic and predicative frameworks [Friedman, 1975; Feferman, 1964; Buss, 1998; Troelstra and van Dalen, 1988; Beeson, 1985]. In this sense, the difficulty of a $2n=p_{1}+p_{2}$ proof exemplifies the general incompleteness of arithmetic when viewed through the lens of constructive realizability.
We close with the central idea: the resolution of future obstructions in mathematics, logic, physics, and computation will not emerge from the isolated treatment of individual problems as dichotomies, but from their systematic integration under a unified arithmetic oracle—an architecture capable of identifying, rather than eliminating, the shared constructive limits that underlie them all:
The last problem is the one whose resolution would render all the others corollaries of its solution.
Acknowledgments
Context
For discussion of $\Phi$-adic” recursion, see arXiv:2510.08934 or Rosko [2025].
Final Remarks
The author welcomes scholarly correspondence and constructive dialogue. No conflicts of interest are declared. This research received no external funding.

### v2

@@
-Manuscript
-Manuscript
-Version One
-ON THE REALIZABILITY OF
-PRIME CONJECTURES
-IN HEYTING ARITHMETIC
-Milan Rosko
-(November 2025)
+ON THE REALIZABILITY OF PRIME CONJECTURES IN HEYTING ARITHMETIC
+Abstract
 We show that no total functional can uniformly transform $\Pi_{1}$ primality into explicit $\Sigma_{1}$ witnesses without violating normalization in $\mathsf{HA}$.
@@
 The formal analysis in constructive logic is followed by heuristic remarks interpreting the results in informational terms.
-Keywords: Proof Theory, Constructive Logic, Realizability;
-Heyting Arithmetic; Goldbach’s Conjecture; Geometric Logic
-MSC: 03F55, 03F03, 11A41, 68Q05, 11U05
-Milan Rosko is from University of Hagen, Germany
-Email: Q1012878 @@ studium.fernuni-hagen.de
-ORCID: 0009-0003-1363-7158
-Licensed under “Deed” ΓΔ  – creativecommons.org/licenses/by/4.0
+Subject: Realizability, Constructive Logic, Primes
+Proof Theory, Recursive Arithmetic, Geometric Logic
 1 Introduction
+Figure 1: Predicting the decomposition of cuboids would effectively provide an oracle.
+Can the semantics of primality (gold) be syntactically separated from the “combinatorial entropy” of decompositions (right)?
+This question frames the present investigation.
 Let mathematics correspond to the greatest rigor after assumption; then logic will be the rigor before assumption.
-We examine the boundary where the two intersect.
+In this spirit, we examine the boundary where the two intersect.
 Classical arithmetic proceeds as though its objects—numbers, primes, and sums—are already well formed, whereas logic asks whether the rules that make such objects definable can themselves support constructive realization.
@@
 It concerns the structure of realizability in arithmetic, not the distribution of prime numbers.
-Figure 1: Predicting the decomposition of cuboids would effectively provide an oracle.
-Can the semantics of primality (gold) be syntactically separated from the “combinatorial entropy” of decompositions (right)?
-This question frames the present investigation.
 The distinction between additive and multiplicative reasoning marks a precise logical frontier.
@@
 Definition 2.2 (Proof Semantics).
-Proofs correspond to explicit constructions (realizers):
+Proofs correspond to realizers:
 […]
@@
 2.2 Thesis
-Example 2.3.
-Recall the Arithmetic Hierarchy of Kleene [1943]. It classifies subsets of $\mathbb{N}$ and arithmetical formulas according to the alternation of quantifiers in their definitions.
+Ansatz 2.3 (BHK Witnessability).
+Throughout, existence is understood in the BHK sense: a statement of the form
+$\forall x\,(\varphi(x)\Rightarrow\exists y\,\psi(x,y))$ (2.3)
+is treated not as a mere classical truth condition, but as a specification of a uniform constructive operator.
+Concretely, a proof in HA of such a sentence is taken to determine a total functional $t$ such that
+$\forall x\,(\varphi(x)\Rightarrow\psi(x,t(x))).$ (2.4)
+Accordingly, we do not regard the possibility that witnesses might exist in principle without being uniformly producible as a semantic resource.
+In this work, arithmetical existence claims are evaluated only insofar as they admit explicit constructive realization under normalization.
+Consider The Teapot of Russell [1952] as an apt anlogy.
+Ansatz 2.4 (Arithmetical Hierarchy).
+Recall the Kleene–Mostowski Hierarchy by Kleene [1943]. It classifies subsets of $\mathbb{N}$ and arithmetical formulas according to the alternation of quantifiers in their definitions.
   1. 1.
@@
     where $\varphi_{0}$ is bounded ($\Delta^{0}_{0}$).
-Thesis 2.4 (Realizability Barrier).
+Thesis 2.5 (Realizability Barrier).
 No total primitive recursive functional $E$ can uniformly transform $\Pi_{1}$ predicates—i.e., elements of $\operatorname{Prime_{\Pi}}$—into explicit $\Sigma_{1}$ witnesses within HA.
@@
 This constraint remains invariant under any external extension, owing to syntactic parsimony.
+Thesis 2.6 (Uniform Realizability Barrier).
+No closed term $t$ in any strongly normalizing realizability calculus adequate for HA (e.g.
+system $\mathcal{T}$ or an equivalent typed $\lambda$-calculus with primitive recursion) can uniformly transform $\Pi_{1}$ evidence into explicit $\Sigma_{1}$ witnesses in the following sense: there is no total realizer $t$ such that, for all $n$,
+$t(n)\;\Vdash\;\Bigl(\Pi_{1}(n)\Rightarrow\exists y\,\Sigma_{1}(n,y)\Bigr)\quad\Longrightarrow\quad\exists y\,\bigl(t(n)=y\wedge\Sigma_{1}(n,y)\bigr).$ (2.9)
+Equivalently, $\Pi_{1}$ predicates do not admit uniform $\Sigma_{1}$-Skolemization within the normalization regime of constructive proof theory.
 2.3 Operators
@@
 $\mathrm{GC}^{*}_{\Pi}:\quad\forall n\Big(\operatorname{Even}(n)\Rightarrow\exists p,q\leq n\,[\operatorname{Prime_{\Pi}}(p)\wedge\operatorname{Prime_{\Pi}}(q)\wedge p+q=n]\Big).$ (2.11)
-This statement combines $\Sigma_{1}$ witnesses $(p,q)$ with $\Pi_{1}$ side-conditions $(\operatorname{Prime_{\Pi}}(p),\operatorname{Prime_{\Pi}}(q))$.
-A constructive proof of (2.8) in HA would yield, via the Curry–Howard Correspondence, a total primitive-recursive operator
-$E:\mathbb{N}\longrightarrow\mathbb{N}^{5},\quad E(n)=\langle p,q,a,b,c\rangle,$ (2.9)
-where $a,b$ realize the predicates $\operatorname{Prime_{\Pi}}(p),\operatorname{Prime_{\Pi}}(q)$, and $c$ realizes $p+q=n$.
-Such an operator would act as a uniform $\Sigma_{1}$ extractor for $\Pi_{1}$ evidence, contradicting the strong normalization of HA proofs, see [Kreisel, 1951].
+Definition 2.9 (Base Calculus).
+Instead of presenting constructive arithmetic solely as a first-order theory, one may work in a typed $\lambda$-calculus $\Lambda_{\mathbb{N}}$ with natural numbers, primitive recursion, and inductive types, whose equational theory corresponds to HA under the Curry–Howard correspondence.
+We call a functional constructively definable if it is represented by a closed term of $\Lambda_{\mathbb{N}}$; equivalently, if it is definable in Gödel’s system $\mathcal{T}$.
+All extracted witness mechanisms discussed below are to be understood in this $\lambda$-definability sense.
+Observation.
+A proof of (2.11) in HA would, under any standard realizability or functional interpretation, induce a uniform closed term $t$ in a strongly normalizing typed $\lambda$-calculus interpreting HA, such that for each even input $n$ the term $t(n)$ computes explicit numerals $\langle p,q\rangle$ satisfying $p+q=n$ together with realizers of the side conditions $\operatorname{Prime_{\Pi}}(p)$ and $\operatorname{Prime_{\Pi}}(q)$.
+In other words, a proof of (2.11) would not merely assert the existence of prime summands, but would entail a uniform constructive mechanism producing them, together with the global $\Pi_{1}$ evidence that they are prime.
+It is precisely this uniformity, rather than any arithmetical property of primes, that contradicts the non-uniformization theorems of constructive proof theory.
+Exposition.
+One might ask: if a uniform witness extractor is unavailable in HA, then in which system could it exist?
+The point is that any such extractor amounts to a global choice principle for $\Pi_{1}$-premised existence, and thus requires principles that exceed normalization-governed constructive arithmetic.
+Typical extensions that admit such uniformization include: (i) classical logic (via $\neg\neg$-elimination and Skolemization), (ii) choice principles for arithmetical predicates, or (iii) non-normalizing computational interpretations (e.g.
+general recursion or bar recursion).
+Our barrier is therefore not contingent on HA syntactically: it persists across all systems whose proof semantics remain strongly normalizing and computationally sound for arithmetic.
 3 Geometric Logic
@@
 $\forall d\in\Delta_{0}\;\exists^{\infty}n\,[\,\mathsf{Comp}(n)\wedge\neg\mathsf{Comp}_{d}(n)\,].$ (3.22)
-Proof Sketch.
 By Diagonalization, see Subsection 5.3. ∎
@@
 As combinatorial complexity increases, local configurations become compressible.
-Proof.
 For any fixed divisor bound $D$, the formula $\mathsf{Comp}_{d}$ can only detect factors $a,b\leq D$.
@@
 $\forall d\,(\text{finite shape-bounded})\ \exists^{\infty}N\,[\mathsf{Comp}(N)\wedge\neg\mathsf{Comp}_{d}(N)].$ (4.29)
-Proof Sketch.
 A divisor a>B(n)a\!>\!B(n) with $a\!\cdot\!b\!=\!n$ escapes any bounded search below $B(n)$.
@@
 with $C$ $\Pi_{1}$-minimal among such sets, then $C(x)\Leftrightarrow\operatorname{Prime_{\Pi}}(x)$.
-Proof Sketch.
 Work in HA. The inclusion $C\subseteq\operatorname{Prime_{\Pi}}$ is given. For the converse, assume $\operatorname{Prime_{\Pi}}(p)$ and $p\notin C$. Consider the $\Pi_{1}$ set
@@
 where $\Phi(e,n)$ states that the $e$-th primitive-recursive operator realizes the $\Sigma_{1}$ witnesses of $\Pi_{1}$ primality required by $\mathrm{GC}^{*}_{\Pi}$.
-Proof Sketch.
 Assume, for contradiction, that
@@
 We simulate classical logic within our constructive framework by inductive reasoning, examining structural correspondences.
-We appeal to the Theory of Inductive Inference, see Solomonoff [1964a, b] and to the Church–Turing Thesis: any realizible arithmetical predicate can be represented by a minimal constructive schema carrying equivalent realization.
+We appeal to the Theory of Inductive Inference, cf.
+Solomonoff [1964a, b] and to the Church–Turing Thesis: any realizible arithmetical predicate can be represented by a minimal constructive schema carrying equivalent realization.
 For primality,
@@
 where $\Theta$ is a bounded arithmetical predicate expressing that $M$ reaches a terminal configuration after $t$ transitions.
-Proof Sketch.
 Unroll the computation as a finite sequence of configurations
@@
 $\frac{\varphi_{n+2}}{\varphi_{n+1}}=1+\frac{1}{\frac{\varphi_{n+1}}{\varphi_{n}}}.$ (5.65)
-Its fixed point is the constant $\Phi$, which quantifies the equilibrium between constructive expansion ($\Sigma_{1}$ search) and reflective contraction ($\Pi_{1}$ verification).
-See Section Context for related discussion.
+Its fixed point is the constant $\Phi$, which quantifies the equilibrium between constructive expansion ($\Sigma_{1}$ search) and reflective contraction ($\Pi_{1}$ verification), see Section Acknowledgments for related discussion.
 Thesis 5.13 ($\Phi$-Adic Equilibrium).
@@
 $\mu_{s}(n)\equiv n^{-s}.$ (5.69)
-Dirichlet addition is ordinary summation, and multiplication acts additively on exponents, see [Apostol, 1976]:
+Dirichlet addition is ordinary summation, and multiplication acts additively on exponents, cf. [Apostol, 1976]:
 $\mu_{s}(ab)=\mu_{s}(a)\mu_{s}(b).$ (5.70)
@@
 $\zeta(s)=\sum_{(\nu_{p})\ \text{finite}}\ \prod_{p\in\mathcal{P}}p^{-\nu_{p}s}=\prod_{p\in\mathcal{P}}\Big(\sum_{k\geq 0}p^{-ks}\Big)=\prod_{p\in\mathcal{P}}\frac{1}{1-p^{-s}}.$ (5.78)
-Proof Sketch.
 Since $\mu_{s}$ is completely multiplicative,
@@
 $\frac{\zeta^{\prime}(s)}{\zeta(s)}=\sum_{p\in\mathcal{P}}\frac{\log p}{p^{s}-1}=\sum_{n\geq 1}\frac{\Lambda(n)}{n^{s}},$ (5.81)
-where $\Lambda$ denotes the Von Mangoldt Function, see [von Mangoldt, 1895].
+where $\Lambda$ denotes the Von Mangoldt Function, cf. [von Mangoldt, 1895].
 Remark.
@@
 $\mathsf{PA}\vdash\mathrm{GC}^{*}_{\Pi}\;\leftrightarrow\;\mathrm{WGC}.$ (6.86)
-Proof Sketch.
 Replace each occurrence of $\operatorname{Prime_{\Pi}}(p)$ in $\mathrm{GC}^{*}_{\Pi}$ by $\exists w\,\operatorname{Cert}(p,w)$ and simplify.
@@
 No such operator $E$ exists.
-Proof Sketch.
 By the Gödel–Blum speedup theorems [Gödel, 1986; Blum, 1967], no arithmetical theory extending $\mathsf{PA}$ can have uniformly bounded proofs for an infinite family of true sentences.
@@
 For example, Quantum computation accelerates instance evaluation but does not transform a given nonconstructive predicate into a realization.
-Logical non-uniformity persists across such computation, see [Abramsky, 2004].
+Logical non-uniformity persists across such computation, cf.
+[Abramsky, 2004].
 Heuristic 7.3 (Constructive Reflection in Quantum Algorithms).
@@
 In this sense, the difficulty of a $2n=p_{1}+p_{2}$ proof exemplifies the general incompleteness of arithmetic when viewed through the lens of constructive realizability.
-We close with the central idea: the resolution of future obstructions in mathematics, logic, physics, and computation will not emerge from the isolated treatment of individual problems as dichotomies, but from their systematic integration under a unified arithmetic oracle—an architecture capable of identifying, rather than eliminating, the shared constructive limits that underlie them all:
+We close with the central idea: the resolution of future obstructions in mathematics, logic, physics, and computation will not emerge from the isolated treatment of individual problems as dichotomies, but from their systematic integration under a general arithmetic oracle—an architecture capable of identifying, rather than eliminating, the shared constructive limits that underlie obstructions:
 The last problem is the one whose resolution would render all the others corollaries of its solution.
 Acknowledgments
-Context
-For discussion of $\Phi$-adic” recursion, see arXiv:2510.08934 or Rosko [2025].
-Final Remarks
-The author welcomes scholarly correspondence and constructive dialogue. No conflicts of interest are declared. This research received no external funding.